OpenStudy (anonymous):

solve..the quaisi..linear equation..... (x^2-y^2-z^2)dz/dx +2xy dz/dy =2xz.... by forming subsidary equations... plz...help!!!!!

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

i have not see quasi linear equations before

OpenStudy (anonymous):

hnm...k..thanks..

OpenStudy (anonymous):

forming the auxiliary equaton...... dx/(x^2-y62-z^2) =dy/(2xy)=dz/2xz........... next?